# DISCOVER: Data-driven Identification of Sub-activities via Clustering and Visualization for Enhanced Recognition of Human Activities in Smart Homes

### v1

@@ 1. Introduction
    Elimination of Pre-Segmentation Constraints: By handling continuous, unsegmented time-series, our framework is applicable to realistic deployment scenarios, and can be easily used in real homes.
@@
    Human-in-the-Loop Annotation with Minimal Effort: Employing an active learning–like strategy, annotators focus on only a small set of representative cluster centroids. Labels then propagate to the entire dataset, drastically reducing manual labeling workload by a factor of over 1,000 (requiring annotation of less than 0.01% of all data points).
  4. (4)
    Open-Source Re-Annotation Tool: We open-source our user-friendly tool that visualizes sensor activations on a 2D house layout, facilitating the labeling of any types of activities.
2. Related Work
To provide context for our work, we review key areas of research that relate to our approach. These include conventional and deep learning-based methods for HAR, which often rely on large labeled datasets. We will then go over advancements in self-supervised and active learning, followed by clustering approaches that aim to reduce the need for manual annotation and be more sample efficient.
2.1. Human Activity Recognition
Human Activity Recognition in smart environments has emerged as a critical area of research, enabling the identification of Activities of Daily Living (ADLs) through ambient sensors (Bouchabou et al., 2021a; van Kasteren et al., 2010; Rafferty et al., 2017; Kientz et al., 2008b). ADLs are particularly valuable for routine detection, which feeds numerous downstream applications, including health monitoring and assistive systems for independent living (Chatting, 2023; Csurka, 2017; Morita et al., 2023). Conventional approaches to HAR rely heavily on supervised learning methods, where large amounts of labeled data is needed. These methods span a range of classic machine-learning techniques, including Random Forests (Sedky et al., 2018), Hidden Markov Models (HMMs) and Support Vector Machines (SVMs) (Cook, 2010). These approaches rely on features hand-crafted by domain experts, and universally assume pre-segmented data, which limits their applicability in real-world scenarios.
Advances in deep learning have expanded the capabilities of HAR systems. Deep Neural Networks (DNNs) (Alaghbari et al., 2022), Long Short-Term Memory networks ((LSTM) (Liciotti et al., 2020), and hybrid architectures such as ELMo combined with BiLSTM (Bouchabou et al., 2021b) have demonstrated improved performance in activity classification tasks. Additionally, graph neural networks (GNNs) have been employed to capture spatial and temporal relationships in sensor data (Li et al., 2019; Zhou et al., 2020; Plötz et al., 2023). Despite these advancements, the reliance on pre-segmented data remains a significant limitation. Some studies have explored segmentation techniques, such as change-point detection algorithms, which identify abrupt changes in sensor data to infer activity boundaries (Aminikhanghahi et al., 2018; Sprint et al., 2020; Jose and Malekian, 2017; Aminikhanghahi and Cook, 2017). However, these methods often rely on heuristics and are not well-suited for capturing fine-grained activities.
These HAR approaches depend on massive labeled datasets and fixed, pre-defined labels that cannot always capture the nuanced behaviors exhibited in real-world environments. To address this challenge, our work introduces a self-supervised framework to process large unlabeled datasets, paired with an active learning approach that allows to discover fine-grained sub-activities with minimal human annotation efforts.
2.2. Self-Supervised Learning
A number of self-supervised learning (SSL) techniques have emerged in computer vision and natural language processing, enabling models to learn generalizable features without relying on massive labeled datasets. This is achieved by designing pretext tasks—such as predicting image rotations (Gidaris et al., 2018) or solving jigsaw puzzles (Noroozi and Favaro, 2016)—that encourage the extraction of robust representations from raw data, which can then be fine-tuned for downstream tasks like image classification and semantic segmentation (Long et al., 2015). A prominent example is SimCLR, a contrastive learning framework that has achieved state-of-the-art results in various domains and has been successfully adapted for time-series data, effectively capturing temporal patterns (Chen et al., 2020; Yang et al., 2022; Shah et al., 2021; Mohsenvand et al., 2020).
In sensor-based HAR, SSL has predominantly been applied in wearable contexts, where the data is homogeneous and continuous, making it well-suited for contrastive and reconstruction-based methods (Haresamudram et al., 2022, 2020). For example, Tang et al. adapt a contrastive learning framework to recognize activities such as walking, sitting, and standing with competitive performance in healthcare applications (Tang et al., 2020). Other methods have enhanced feature extraction by incorporating custom transformer encoders (Khaertdinov et al., 2021) and clustering-based negative selection (Wang et al., 2023) into their frameworks.
However, smart home HAR presents a different challenge: the data is heterogeneous, combining discrete events with continuous sensor streams, which complicates the direct application of traditional SSL methods. Chen et al. (Chen et al., 2020) applied SSL to ambient sensor data by developing a SimCLR-based model that trains an encoder to capture both spatial and temporal dependencies, extracting robust features for downstream HAR tasks and achieving superior performance in semi-supervised and transfer learning scenarios. The main goal of their approach is to predict the original coarse labels. We, however, go beyond that paradigm, aiming to uncover and validate semantically meaningful, fine-grained sub-activities within the complex domain of smart homes.
2.3. Active Learning
The annotation of sensor data for HAR is a resource-intensive process, requiring significant time and expertise (Hiremath and Plötz, 2020; Zhai et al., 2019). Active learning has emerged as a promising approach to reduce the annotation burden by selecting the most informative samples for labeling (Settles, 2009). This human-in-the-loop paradigm aims to minimize the number of labeled samples required while maintaining model performance. A key challenge in active learning is determining an optimal sampling strategy, oftentimes having to choose between pool-based and stream-based approaches (Hossain et al., 2017; Adaimi and Thomaz, 2019; Jones et al., 2014). In contrast, our approach utilizes a clustering-based active learning framework that identifies representative sub-activities for annotation. This minimizes human intervention while retaining the flexibility to dynamically adapt to novel data sources.
2.4. Clustering
Unsupervised clustering methods have been employed in HAR to group unlabeled sensor data into meaningful activity clusters (Ariza Colpas et al., 2020). Classic approaches, such as Lloyd’s clustering algorithm combined with k-Nearest Neighbors (Fahad et al., 2014) and DBSCAN (Hoque and Stankovic, 2012), have been used to identify patterns in sensor data. More recently, the SCAN framework (Gansbeke et al., 2020) has been adapted for HAR, enabling the clustering of unlabeled movement data from wearable sensors (Ahmed et al., 2022). These methods leverage large-scale sensor data to enhance activity recognition, addressing the challenges associated with obtaining annotated datasets. However, existing ”cluster then classify” approaches still rely on downstream labels to match and often lack the ability to discover fine-grained activities.
Recent work by Bock et al. (Bock et al., 2024) introduced a clustering-based annotation pipeline that reduces the need for human labeling by leveraging vision-based models for weak annotation in HAR datasets, demonstrating that clustering methods can match fully supervised classifiers while minimizing manual effort. Inspired by successes in the wearable domain, we propose a novel approach tailored for ambient sensor data in smart homes that addresses the challenges of heterogeneous sensor streams while discovering fine-grained sub-activities rather than merely mapping back to the original labels.
3. DISCOVER: Data-driven Identification of Sub-activities via Clustering and Visualization for Enhanced Recognition
Figure 1. Overview of DISCOVER, a self-supervised system designed to discover fine-grained human activities from unlabeled sensor data without relying on pre-segmentation, consisting of two main stages: clustering and labeling. After (0) slicing the raw data into continuous sliding windows of sensor activations without assuming any pre-segmentation, we (1) train a BERT model with mask language modeling task to encode the windows in an embedding space. We then use these embeddings to identify similar activity windows and (2) fine-tune a clustering model using SCAN loss, which results in assigning all data points to $k$ clusters. We then (3) sample a handful of windows closest to each cluster centroid, replay them on 2D house layouts using our custom built visualization tool, and send these samples to a group of experts for annotation. With minimal labeling effort, we obtain custom granular activity labels for each cluster centroid, and (4) propagate them to the rest of the data points in each cluster. These custom labels are then (5) applied to the original dataset and can later be used for a set of specialized downstream tasks.
To address the challenges of human activity recognition in smart homes, we introduce DISCOVER, a data-driven approach that discovers fine-grained human activities from unlabeled sensor data without relying on pre-segmentation. DISCOVER combines unsupervised feature extraction with an interactive annotation tool, enabling more granular and personalized activity labels.
We consider the problem of human activity recognition in smart home environments. Given a time-series dataset of sensor readings collected over a period of time, divided into a collection of $n$ sliding windows $\{W_{1},W_{2},\dots,W_{n}\}$ of size $l$, our goal is to assign each of these windows to an activity label. Formally, we seek to learn a mapping $f:W_{i}\mapsto c_{k},$ where $c_{k}$ represents the activity cluster for window $W_{i}$. A natural challenge in this setting is the lack of ground truth labels for the time windows.
Fig. 1 gives an overview of DISCOVER, which consists of two main stages: clustering and labeling. The approach starts by pre-processing sensor data into sliding windows in step (0). We then pursue a active learning sub-activity discovery through the following five steps:
  1. (1)
    Encoder Pre-Training: Create an embedding representation of sensor readings by training a BERT transformer model with a self-supervised, mask language modeling task. Use these embeddings to create a set of $m$ nearest neighbors for each window $W_{i}$.
  2. (2)
    Clustering Model Fine-Tuning: Fine-tune the pre-trained BERT model using the SCAN loss function (Gansbeke et al., 2020), which partitions all windows $W_{i}$ into $k$ distinct clusters, assigning each $W_{i}$ to one of them.
@@ 2.4. Clustering
    Re-Annotation of the Original Time-Series Data: Apply these cluster activity labels to the original time-series data.
@@ 3.1. Encoder Pre-Training
Figure 2. DISCOVER approach – model training pipeline, consisting of (1) Encoder Pre-Training, and (2) Clustering Model Fine-Tuning. DISCOVER first trains a BERT model using a Masked Language Modeling head in (1) to obtain initial embeddings for each window $W_{i}$. In step 2(a) it uses these embeddings to identify similar activity windows and pairs them together as a a new training set. It then continues training the pre-trained BERT base model with a SCAN loss (2(b)). In the end, the trained SCAN model assigns a cluster $c_{k}$ to each input sequence $W_{i}$ (2(c)).
The first step is to learn an informative representation for raw sensor sequences. In this subsection, we describe how we adapt a BERT model through a masked language modeling (MLM) objective to capture context-rich representations of ambient sensor data without any labels.
We build on the approach from (Hiremath et al., 2022), which employs a pre-trained BERT Transformer model to derive embeddings for sensor sequences by feeding in new, domain-specific tokens and training for several epochs. However, unlike (Hiremath et al., 2022), our pipeline does not rely on pre-segmented data, staying faithful to a more realistic, fully unsupervised scenario. Fig. 2(1) shows a diagram of this process. First, we load the original BERT model that was pre-trained on text corpora (BooksCorpus and English Wikipedia) (Devlin et al., 2019) from HuggingFace (Wolf et al., 2020) and extend the original vocabulary of size 30,522 WordPiece tokens by adding sensors and their readings as new tokens to the original vocabulary (which evaluates to roughly 100 new tokens per dataset). For example, sensor labeled M1 and its output signal ON are concatenated into a new token M1_ON, assigned a new token ID, and initialized with random 768-dimensional embedding. Once all training input sequences are processed and encoded, we train BERT model using a self-supervised MLM objective, defined below.
@@
The training loss $\mathcal{L}_{MLM}$ is defined as:
$\mathcal{L}_{MLM}=-\frac{1}{|M|}\sum_{x\in M}\log P(d_{x}\mid\hat{W_{i}})$
where $P(d_{x}\mid\hat{W_{i}})$ is the probability that $x$-th masked token $d$ is correctly predicted by the model.
This loss function incorporates the context from the surrounding, unmasked tokens (i.e., sensor readings) in each sequence. By introducing new tokens for sensor readings (e.g., M1_ON) into the vocabulary and initializing their embeddings randomly, the model is able to integrate sensor-specific information into the learned representations during pre-training. Once the training is done, we obtain an embedding for each input sequence $W_{i}$ by removing the MLM head and extracting the 768-dimensional vector of a special [CLS] token. This [CLS] (”classification”) token is specifically designed to represent the entire sequence in a single embedding, a common practice in BERT-based architectures to obtain a fixed-length representation of any input sequence (Devlin et al., 2019).
@@ 3.2. Fine-Tuning the Clustering Model
Once we obtain an embedding for each window $W_{i}$ in the training data, we use these embeddings to measure the similarity between individual sequence vectors, build a ”neighbors” dataset, and train a clustering model using this data.
First, for every input sequence we want to find and store a small set of examples that are most similar to it in the BERT embedding space (Fig. 2 (2a)). Let $\text{sim}(W_{q},W_{r})$ represent the cosine similarity between two embedding vectors $W_{q}$ and $W_{r}$, defined as:
@@
For every input sequence $W_{i}$, we identify its $m$-nearest neighbors using this cosine similarity. Following (Gansbeke et al., 2020), we set $m$ equal to 20. These neighbors will be used as training examples in the following step where the SCAN loss function will be forcing them to be in the same cluster $c_{k}$
Then, we further fine-tune the re-trained BERT model from the previous step by redefining its training objective (Fig. 2 (2b)). To achieve that, we replace the MLM training head and its cross-entropy loss function with a clustering head paired with the SCAN loss (Gansbeke et al., 2020). SCAN loss consists of two components, an instance-level contrastive loss and a cluster-level entropy loss, defined as follows:
  • Instance-Level Contrastive Loss: Encourages every input sequence and its neighbors to belong to the same cluster. For a sequence $W_{i}$, let $M(W_{i})$ denote its set of $m$-nearest neighbors. The contrastive loss $\mathcal{L}_{\text{contrastive}}$ is defined as:
    $\mathcal{L}_{\text{contrastive}}=-\frac{1}{n}\sum_{W_{i}}\frac{1}{|M(W_{i})|}\sum_{W_{j}\in M(W_{i})}\log P(c_{i}=c_{j})$
@@
    where $\lambda$ controls the weight of the entropy loss. We use the default value of $\lambda=2$ (Gansbeke et al., 2020).
@@ 3.3. Cluster Centroid Annotation
Figure 3. DISCOVER custom-built interactive in-browser annotation tool for reviewing sensor activation sequences. The tool displays a 2D house layout, allowing annotators to replay sequences temporally, observe contextual and spatial details, and assign labels via a drop-down menu. The example above shows a sequence of sensor activations following a resident walking to the guest bathroom, that an annotator can label as ”Guest Bathroom: Walking In” from the drop-down menu.
Given the set of clusters learned in the previous stage, the next goal is to interpret and contextualize them by assigning meaningful labels. Since the model has so far operated without any supervision, no ready-made labels exist to inform downstream analyses or practical applications. By annotating a small number of representative samples from each cluster, we can translate the resulting clusters into actionable units—such as identifying specific sub-activities or understanding when and where a resident might be performing a certain activity. This process not only validates the clusters internally but also bridges the gap between the automatic discovery of patterns and real-world interpretability.
To minimize the number of sequences we need to review, we will only use $N$ samples from each cluster that have the highest probability of belonging to that cluster. The probability $P(c_{i}=k)$ is taken directly from the SCAN model output. In step (3) in Fig. 1, we select $N\times k$ samples and review them manually using the custom tool that we developed, shown in Fig. 3. This interactive in-browser annotation tool shows a detailed layout of a given house floor plan and allows to playback the sensor activation sequences temporally, and assign a label from a drop-down menu. This tool lets the annotator see both the contextual information about that sequence (e.g., location of the activity and its surrounding areas), and the temporal dimension (e.g., whether the activity was static, or if it triggered multiple sensor activations throughout the house). For example, Fig. 3 displays a sequence of sensor activations representing a resident walking through the house to the guest bathroom. The annotator can label this sequence as ”Guest Bathroom: Walking In.” DISCOVER visualization tool simplifies and speeds up the annotation process and allows for faster iterations.
@@ 3.5. Re-Annotation of the Original Time-Series Data
4. Experimental Evaluation: Data
Annotated sample sets are essential for HAR in smart homes, but they are often difficult to obtain. In this work, we present a method that assists human annotators in generating labeled datasets, potentially guiding data recording as well. In the following sections, we first provide an overview of the datasets used for our experimental validation and evaluation in Section 4.1 and summarize the relevant pre-processing applied to the data in Section 4.2. Then, we validate the quality and relevance of the outputs produced during the clustering stage (see Fig. 1) in Section 5 – establishing the effectiveness of our approach. Finally, we evaluate the quality of the Labeling stage of the pipeline and explore the quality and relevance of our finer-grained annotation scheme in Section 6.
@@ 4.1. Datasets
[…]
Table 1. CASAS Datasets: Label Distribution Summary for Milan, Aruba, and Cairo
Our experimental analysis uses publicly available labeled datasets from the Center of Advanced Studies in Adaptive Systems (CASAS) at the Washington State University (Cook et al., 2009). The CASAS testbed consists of multiple household datasets, each containing motion (M), door (D), and temperature (T) sensor streams collected over a period of time that stretches multiple months, with associated activity labels. In this work, we perform analysis on three households that are widely used in prior HAR work (Demirel et al., 2018; Thukral et al., 2024):
@@
Each CASAS dataset includes assigned activity labels, capturing a range of household activities, such as cooking, relaxing, sleeping. While the labels broadly overlap across datasets, subtle discrepancies exist: for example, “guest bathroom activity” in Milan does not appear in Aruba or Cairo, as there are no guest bathroom sensors in those households. Following (Demirel et al., 2018; Thukral et al., 2024), we resolve these discrepancies by mapping each dataset’s labels to a unified set of standard categories (e.g., Relax, Cook, Sleep), as detailed in (Thukral et al., 2024).
The activity label distribution is far from uniform, both within each dataset and between the households. For instance, in Milan, approximately 29% of sensor readings correspond to cooking activities, whereas only 3% are related to work. There’s also significant inter-house variability: in Milan, Cook accounts for 29% of all data points, compared to only 17% in Aruba and effectively none in Cairo. All three datasets also incorporate an Other category, that serves as a catch-all for both rare activities like grooming and meditating and unlabeled periods of sensor readings that do not fit any specified category. As shown in Table 1, the prevalence of Other ranges widely, from 33% of data points in Milan, to 53% in Aruba and 77% in Cairo. In the context of DISCOVER pipeline, however, Other is less useful as a standalone activity category, since even seemingly trivial movements such as walking between rooms could constitute a valid sub-activity. This distinction will be important when we benchmark DISCOVER against fully-supervised state-of-the-art approaches.

### v3

@@ 2.5. Clustering Pipelines
    Encoder Pre-Training: Create an embedding representation of sensor readings by training a transformer model with a self-supervised masked language modeling objective. Transformers have been shown to be effective for modeling long-range structure in event-driven sensor streams (You et al., 2025b), and we treat sequences of ambient sensor events analogously to sentences. This enables the model to learn context-aware representations that capture temporal dependencies and co-occurrence structure in daily routines. We then use these embeddings to create a set of $m$ nearest neighbors for each window $W_{i}$.
@@
    Clustering Model Fine-Tuning: Fine-tune the pre-trained encoder using SCAN loss (Gansbeke et al., 2020), which encourages consistency among nearest-neighbor embeddings and partitions all windows $W_{i}$ into $k$ clusters. This serves as our primary clustering approach for discovering data-driven Patterns of Daily Living.
@@
    Re-Annotation of the Original Time-Series Data: Apply these cluster labels to the original time-series data.
The following sections describe the technical details of each stage of DISCOVER, beginning with self-supervised representation learning, followed by clustering-based pattern discovery, and concluding with human-in-the-loop annotation and label propagation.
@@ 3.1. Encoder Pre-Training
Figure 2. DISCOVER model training pipeline, consisting of (1) Encoder Pre-Training, and (2) Clustering Model Fine-Tuning. DISCOVER first trains a Transformer model using a Masked Language Modeling head in (1) to obtain semantic embeddings for each window $W_{i}$. In step 2(a) it uses these embeddings to identify similar activity windows and pairs them together as a a new training set. It then continues training the fine-tuning the Transformer model with a SCAN loss (2(b)). In the end, the trained SCAN model assigns a cluster $c_{k}$ to each input sequence $W_{i}$ (2(c)).
The first step is to learn an informative representation from raw sensor sequences. In this subsection, we describe how we train a Transformer sequence model using a masked language modeling (MLM) objective to capture context-rich representations of ambient sensor data without requiring any labels.
We build on the approach from Hiremath et al. (Hiremath et al., 2022), which employs a pre-trained BERT Transformer model to derive embeddings for sensor sequences by feeding in new, domain-specific tokens and training for several epochs. In addition to the advantages demonstrated in (Hiremath et al., 2022), we also note that recent work in the ambient-sensor HAR community has shown that Transformer-based architectures outperform recurrent models such as BiLSTMs. In particular, You et al. (You et al., 2025b) demonstrate that Transformers provide better modeling of long-range temporal dependencies, handle sparse event-driven sequences more effectively than recurrent architectures, and offer improved generalization across heterogeneous home layouts by learning richer contextual structure. Unlike the approach proposed by Hiremath et al. in (Hiremath et al., 2022), our pipeline does not rely on pre-segmented data, staying faithful to a more realistic, fully unsupervised scenario. Fig. 2 (1) shows a diagram of this process. First, we load the original BERT model that was pre-trained on text corpora (BooksCorpus and English Wikipedia) (Devlin et al., 2019) from HuggingFace (Wolf et al., 2020) and extend the original vocabulary of size 30,522 WordPiece tokens by adding sensors and their readings as new tokens to the original vocabulary (which evaluates to roughly 100 new tokens per dataset). For example, sensor labeled M1 and its output signal ON are concatenated into a new token M1_ON, assigned a new token ID, and initialized with random 768-dimensional embedding. Once all training input sequences are processed and encoded, we train BERT model using a self-supervised MLM objective, defined below.
@@
The training loss $\mathcal{L}_{MLM}$, which is essentially a cross-entropy over the masked tokens, is defined as:
[…]
This loss function incorporates the context from the surrounding, unmasked tokens (i.e., sensor readings) in each sequence. By introducing new tokens for sensor readings (e.g., M1_ON) into the vocabulary and initializing their embeddings randomly, the model is able to integrate sensor-specific information into the learned representations during pre-training. Once the training is done, we obtain an embedding for each input sequence $W_{i}$ by removing the MLM head and extracting the 768-dimensional vector of a special [CLS] token. This [CLS] (“classification”) token is specifically designed to represent the entire sequence in a single embedding, a common practice in BERT-based architectures to obtain a fixed-length representation of any input sequence (Devlin et al., 2019).
@@ 3.2. Fine-Tuning the Clustering Model
Once we obtain an embedding for each window $W_{i}$ in the training data, we use these embeddings to measure the similarity between individual sequence vectors, build a “neighbors” dataset, and train a clustering model using this data. First, for every input sequence we want to find and store a small set of examples that are most similar to it in the BERT embedding space (Fig. 2 (2a)). Let $\text{sim}(W_{q},W_{r})$ represent the cosine similarity between two embedding vectors $W_{q}$ and $W_{r}$, defined as:
@@
For every input sequence $W_{i}$, we identify its $h$-nearest neighbors using this cosine similarity. Following the original SCAN paper (Gansbeke et al., 2020), we set $h$ equal to 20. These neighbors will be used as training examples in the following step where the SCAN loss will be forcing them to be in the same cluster $c_{k}$
We take the re-trained BERT model from the previous step and fine-tune it by redefining its training objective (Fig. 2 (2b)). To achieve that, we replace the MLM training head and its cross-entropy loss function with a clustering head paired with the SCAN loss (Gansbeke et al., 2020). SCAN loss consists of two components, an instance-level contrastive loss and a cluster-level entropy loss, defined as follows:
  • Instance-Level Contrastive Loss: Encourages every input sequence and its neighbors to belong to the same cluster. For a sequence $W_{i}$, let $H(W_{i})$ denote its set of $h$-nearest neighbors. The contrastive loss $\mathcal{L}_{\text{contrastive}}$ is defined as:
    $\mathcal{L}_{\text{contrastive}}=-\frac{1}{n}\sum_{W_{i}}\frac{1}{|H(W_{i})|}\sum_{W_{j}\in H(W_{i})}\log P(c_{i}=c_{j})$
@@
    where $\lambda$ controls the weight of the entropy loss. Following (Gansbeke et al., 2020), we set $\lambda=2$.
@@ 3.3. Cluster Centroid Annotation
Figure 3. DISCOVER custom-built interactive in-browser annotation tool for reviewing sensor activation sequences (Karpekov et al., 2025) The tool displays a 2D house layout, allowing annotators to replay sequences temporally, observe contextual and spatial details, and assign labels via a drop-down menu. The example above shows a sequence of sensor activations following a resident walking to the guest bathroom. The annotator decides to label this sample as “Guest Bathroom: Walking In”.
Given the set of clusters learned in the previous stage, the next goal is to interpret and contextualize them by assigning meaningful labels. Since the model has so far operated without any supervision, no ready-made labels exist to inform downstream analyses or practical applications. By annotating a small number of representative samples from each cluster, we can translate the resulting clusters into actionable units—such as identifying specific patterns of daily living or understanding when and where a resident might be performing a certain activity. This process not only validates the clusters internally but also bridges the gap between the automatic discovery of patterns and real-world interpretability.
To minimize the number of sequences we need to review, we will only use $m$ samples from each cluster that have the highest probability of belonging to that cluster. The probability $P(c_{i}=k)$ is taken directly from the SCAN model output. In step (3) in Fig. 1, we select $m\times k$ samples and review them manually using a custom visualization tool, shown in Fig. 3 (Karpekov et al., 2025). This interactive in-browser annotation tool shows a detailed layout of a given house floor plan and allows to playback the sensor activation sequences temporally, and to assign a label from a drop-down menu. This tool lets the annotator see both the contextual information about that sequence (e.g., location of the activity and its surrounding areas), and the temporal dimension (e.g., whether the activity was static, or if it triggered multiple sensor activations throughout the house). For example, Fig. 3 displays a sequence of sensor activations representing a resident walking through the house to the guest bathroom. The annotator can label this sequence as “Guest Bathroom: Walking In”. DISCOVER visualization tool simplifies and speeds up the annotation process and allows for faster iterations.
@@ 3.5. Re-Annotation of the Original Time-Series Data
4. Data Pre-processing and Experimental Setup
In this section, we provide an overview of the datasets used for our experimental validation and summarize pre-processing steps applied to the data, and describe our labeling tool.
@@ 4.1. Datasets
Label Milan Aruba Cairo Cook 29% 17% - Eat 1% 1% 15% Relax 16% 22% - Work 3% 2% 1% Sleep 5% 4% 2% Bathing 6% - - Bed_to_toilet 0.28% - 0.23% Take_medicine 0.13% - 0.28% Leave_Home 1.01% 0.11% 0.12% Other 5% 0.03% 4% No Label 33% 53% 77% TOTAL # sensor readings 433,665 1,716,039 726,534 Total # Unique days of data 72 220 58 Num labels 9 6 6
Table 1. CASAS Datasets (Cook et al., 2009): Label Distribution Summary for Milan, Aruba, and Cairo
Our experimental analysis uses publicly available datasets from the Center of Advanced Studies in Adaptive Systems (CASAS) at the Washington State University (Cook et al., 2009). The CASAS testbed consists of multiple household datasets, each containing motion (M), door (D), and temperature (T) sensor streams collected over a period of time that stretches multiple months, with associated activity labels. In this work, we perform analysis on three households that are widely used in prior HAR work (e.g., (Liciotti et al., 2020; Thukral et al., 2024)):
@@
Each CASAS dataset includes assigned activity labels, capturing a range of household activities, such as cooking, relaxing, sleeping. While the labels broadly overlap across datasets, some discrepancies exist: for example, “guest bathroom activity” in Milan does not appear in Aruba or Cairo, as there are no guest bathroom sensors in those households. When comparing our approach to CASAS labels, we resolve discrepancies by mapping each dataset’s labels to a unified set of standard categories (e.g., Relax, Cook, Sleep), following the method established in prior work (Liciotti et al., 2020; Thukral et al., 2024).
Supporting our motivation for data-driven discovery of patterns of daily living instead of a fixed set of activity labels, we note that a significant proportion of CASAS data has no activity label (No Label), ranging from 33% of data points in Milan, to 53% in Aruba and 77% in Cairo. Prior HAR techniques that rely on pre-segmentation excluded this data from their analyses (Thukral et al., 2024; Demirel et al., 2018; Liciotti et al., 2020; Bouchabou et al., 2021b, c; You et al., 2025a; Gao et al., 2024; Dhekane and Ploetz, 2025a), thereby eliminating 33-77% of human behavior data from their analysis. As our results will demonstrate, DISCOVER enables us to cluster and obtain labels for all data, ensuring a more complete analysis of underlying human behavior patterns.